OpenStudy (anonymous):

Write the equation y - 4 = -2(x - 5) in slope-intercept form. Answer y = -2x + 14 y = -2x + 5 y = -2x + 9 y = -2x + 6

OpenStudy (solomonzelman):

expand the parenthesis on the right hand side add 4 to both sides.

OpenStudy (solomonzelman):

y - 4 = -2(x - 5) y - 4 = -2x + 10 y - 4 + 4 = -2x + 10 + 4 y = -2x + 14